OpenStudy (anonymous):

ok So what are the values of the funtion y=-2 - 4 for x=0, 1, 2, and 3

OpenStudy (anonymous):

is there a x in the function?

OpenStudy (anonymous):

x is the -2x

OpenStudy (anonymous):

-2x -4?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

mk so plug in each one first x= 0 so -2 (0) - 4 = ?

OpenStudy (anonymous):

when x=1 it would be -2 (1) -4 =? and so on

OpenStudy (anonymous):

like -2 (0) = 0 so 0- 4 = -4 for x=0

OpenStudy (anonymous):

So x is 0?

OpenStudy (anonymous):

nononononono lemme write this out for the first 2

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

I get it